The reaction A(aq) ---> B(aq) + C(aq) is a first order reaction. The half-life of A(aq) is 86.6 s at 25.0oC and its half-life is 66.2 s at 45.0oC. What is its half-life (in s) at 65.0oC?
Answer #1

Answer:-

This question is solved by using the simple concept of Arrhenius equation which involves the relationship rate constant and the activation energy.

The answer is given in the image,

Answersa Alacgy Beq) +C64) First order reaction K = 0.693 - ttg kr 30-693 86.5 = 8.00 X10-352 kg=0.693 66.2 = 0.0105 52 Ti =